OpenStudy (anonymous):

Use point-slope to write an equation of a line with slope of -7/9 and passing through the point (5,3). Write final equation in slope-intercept form

jigglypuff314 (jigglypuff314):

point-slope form: y - y1 = m(x - x1) when given slope = m and point (x1, y1) so plug in m = -7/9 and x1 = 3 and y1 = 3 then once you get your equation solve so that you get just y on one side which would then be slope-intercept form